# Biased Backpressure Routing Using Link Features and Graph Neural Networks

### v1

@@ Abstract
Index Terms:
Backpressure, MaxWeight scheduling, shortest path, queueing networks, last packet problem, graph neural networks.
publicationid: pubid: $0$–$00$ 2024 IEEE
I Introduction
@@
There are two categories of low-overhead BP variations for latency improvement: 1) Biased BP adds pre-defined queue-agnostic biases, e.g., (functions of) shortest path distance [12, 13, 15, 1, 2] to the backlog metric. Shortest path-biased BP (SP-BP) can mitigate the drawbacks of slow startup and random walk, while maintaining throughput optimality [12, 13, 15] at a low cost, i.e., a one-time communication overhead for bias computation. As exemplified in Fig. 1(b), two major routes around the empty central area can be quickly established by SP-BP [12, 13]. 2) Delay-based BP employ delay-based backlog metrics rather than queue length [19, 16, 21], in order to address the last packet problem and to stabilize the queues in the network. Importantly, the aforementioned approaches do not require additional communication overhead at every time step, unlike those based on queue state information (QSI) of the network [16, 17] or virtual queues [20, 18] (see Section II-A for a more detailed survey). However, to get the most out of these approaches, careful parameter tuning is required, typically done via trial-and-error.
@@
II Related Work
II-A Backpressure Routing
@@
II-B GNNs for Wireless Resource Allocation
@@
III System Model and Backpressure Algorithm
@@ III-B Backpressure Algorithm
where $U_{i}^{(c)}(t)$ is the backlog metric, whose relationship with queue lengths is detailed in Section III-C. For shorter notation, we define the backpressure of commodity $c$ on the directed link ($\overrightarrow{i,j}$) as $U_{ij}^{(c)}(t)=U_{i}^{(c)}(t)-U_{j}^{(c)}(t)$. In step 2, the maximum backpressure of ($\overrightarrow{i,j}$) is found as:
@@
III-C (Biased) Backlog Metrics
@@
IV Link Duty Cycle Prediction with GNNs
Existing SP-BP algorithms are based on the hop distance (minimal hop count) between two nodes in the networks, which implicitly assumes that every link contributes equally to the end-to-end delay of routing. This, however, ignores the fact that the average delay of each link is different due to its network locality and link scheduling policy. In MaxWeight scheduling, the likelihood of a link being scheduled depends on its local conflict or interference topology and the network traffic load. We propose using a graph convolutional neural network (GCNN) to predict the expected delay on each link, which can serve to design a shortest path bias between node $i$ and the destination node $c$ in (6) better than the shortest hop distance. The GNN and its training setup are illustrated in the upper right blocks of Fig. 2, while the optimal scaling of link weights is discussed in Section V-A. Since our focus is on model-based framework that enables the applications of GNNs for enhancing SP-BP, we select GCNN for its simplicity and good generalizability, leaving the search of optimal GNN architectures and hyperparameters for the future work.
IV-A Graph Convolutional Neural Networks
@@
IV-B Delay-Aware Edge Weights
Depending on the availability of the long-term link rate $r_{e}$ and link duty cycle $x_{e}$ for any link $e\in{\mathcal{E}}$, we consider the weight of each link for finding the shortest path distance between any pair of nodes to be defined in three possible ways: $\delta_{e}=1/x_{e}$, $\delta_{e}=1/r_{e}$, or $\delta_{e}=\bar{r}/(x_{e}r_{e})$, where $\bar{r}$ is the network-wide average link rate defined in Section III-A. In addition, these link weights can be optimally scaled by $\bar{r}/\delta_{\min}$ based on the principle in Section V-A. The computation of vector $\bm{\delta}=\left[\delta_{e}|e\in{\mathcal{E}}\right]$ is denoted as a function $\bm{\delta}=f_{w}({\mathbf{x}},{\mathbf{r}},\bar{r})$. By setting the edge weights of the connectivity graph ${\mathcal{G}}^{n}$ as $\bm{\delta}$, bias $B_{i}^{(c)}$ is set as the weighted shortest path distance between nodes $i$ and $c$ on ${\mathcal{G}}^{n}$. This distance can be computed with distributed algorithms, such as weighted APSP in general or weighted SSSP when a new node joins the network.
IV-C Complexity
@@ IV-D Training
Intuitively, by minimizing the loss in (9), we are choosing parameters $\bm{\omega}$ such that the softmax output of our GCNN ${\mathbf{X}}^{L}$ is close to predicting the fraction of time that each link is scheduled. We do not seek to learn this complex function for any specific topology, but rather, we want to minimize the average error over instances drawn from $\Omega$. In practice, we approximate the expected values in (9) with the corresponding empirical averages. Indeed, with the collected experience tuples, we update the parameters $\bm{\omega}$ of the GCNN after each training instance, through batch training with random memory sampling, employing the Adam optimizer. Notice that, once trained, the GCNN can be used to compute the delay-aware bias for previously unseen topologies without any retraining. As long as the new topology, arrival rates, and link rates are similar to those observed in $\Omega$ during training, we illustrate in Section VII that the trained GCNN generalizes to new test instances not observed during training.
V Improvements to Biased Backpressure
@@
V-A Optimal Scaling of Edge Weights
To the best of our knowledge, the constant scale of the edge weights $\delta_{e},e\in{\mathcal{E}}$ for computing the shortest path biases in SP-BP schemes are often selected via trail-and-error, such as a constant $\delta_{e}$ for all $e\in{\mathcal{E}}$ in EDR [12] and $\delta_{e}=\bar{r}/(x_{e}r_{e})$ proposed in Section IV-B. Our goal is to provide a principled method of optimization for such scaling parameter(s).
@@ V-A Optimal Scaling of Edge Weights
Proof.
In queue length-based SP-BP, the backpressure on link $(\overrightarrow{i,j})$ is $U_{ij}^{(c)}\!(t)=Q_{ij}^{(c)}\!(t)+\big{(}B_{i}^{(c)}\!-\!B_{j}^{(c)}\big{)}=q+\delta_{e}$. Since link $(\overrightarrow{i,j})$ is scheduled at time $t$, we have that $U_{ij}^{(c)}\!(t+1)=\delta_{e}-q$. To avoid the direction of the backpressure on link $(i,j)$ being reversed at time $t+1$, i.e., $U_{ij}^{(c)}\!(t\!+\!1)<0$, we need to set $\delta_{e}\geq\bar{r}$. ∎
@@
Although the above discussion is based on the CFLP setting with all link rates equal to the average $\bar{r}$, in Section VII, we demonstrate experimentally the optimality of the above choice under heterogeneous link rates and general traffic settings.
V-B Bias Maintenance
@@
V-C Delay-aware backlog metric: expQ
@@
where $\epsilon\geq 0$ is a small constant parameter, and it follows that $g\big{(}{\mathcal{Q}}_{i}^{(c)}(t)\big{)}=Q_{i}^{(c)}(t)$ for $\epsilon=0$ and that $g\left(\varnothing\right)=0$ for an empty queue $\varnothing$. With this definition, it is easy to find that $g\big{(}{\mathcal{Q}}_{i}^{(c)}(t)\big{)}\geq Q_{i}^{(c)}(t)$. Further restrictions on $\epsilon$ can be found in (25) and the following discussion in Section VI-B.
@@ VI Throughput Optimality
VI-A Queue length-based SP-BP scheme
@@
By squaring the queue update in (15), and using the fact that $(\max[q-o,b]+p)^{2}\leq q^{2}+o^{2}+p^{2}+b^{2}+2bp+2q(p-o)$ for $q\geq 0,o\geq 0,b\geq 0,p\geq 0$ (see Appendix A), we have:
@@
Next, we show that the biased BP described in Section III-B is designed to choose a feasible policy ${\mathbf{M}}(t)=\big{(}\mu_{ij}^{(c)}(t)\big{)}_{i,j,c\in{\mathcal{V}}}$ under current network state to minimize the bound of the Lyapunov drift, i.e., the right-hand side of (16). Since $W$ and $\lambda_{i}^{(c)}$ are constants, this amounts to minimizing the following:
@@
Since $W\!<\!\infty$, $0\!\leq\!B_{i}^{(c)}\!<\!\infty$, the condition of strong stability in (12) immediately follows by moving $\sum_{i,c\in{\mathcal{V}}}B_{i}^{(c)}$ from the left hand side of (21) to the right hand side. ∎
VI-B SP-BP scheme with expQ
Theorem 2.
There exists a constant $\epsilon\geq 0$ such that SP-BP with delay-based backlog metric expQ defined in Section V-C, denoted as $E_{i}^{(c)}\!(t)=g\!\left(\!{\mathcal{Q}}_{i}^{(c)}(t)\!\right)$, meets the condition of strong stability in (12) within a reduced network capacity region.
Proof.
@@ VI-B SP-BP scheme with expQ
We apply the proof for Theorem 1 by substituting (15) with (22), and replacing $U_{i}^{(c)}(t)$ by $\tilde{U}_{i}^{(c)}(t)$ and $A_{i}^{(c)}\!(t)$ by $A_{i}^{(c)}\!(t)+b$, thus obtain the policy of SP-BP with expQ as
@@
With (24), (25) and ${E}_{i}^{(c)}\!(t)\geq{Q}_{i}^{(c)}\!(t)$, all queues are strongly stable within the network capacity region shrank by $b$. ∎
@@
Simulations in [13, Ch 4.9] and Section VII-C show that for BP with queue length-based backlog metrics, the inclusion of non-negative queue-agnostic biases does not shrink the network capacity region of the basic BP. However, as illustrated in Section VII-C, the application of delay-based backlog metrics, including expQ, will reduce the network capacity of BP.
VII Numerical experiments
@@
To illustrate the effectiveness of our proposed enhancements to the SP-BP schemes in improving routing performance, we evaluate a series of low-overhead BP routing algorithms in simulated wireless multi-hop networks. Each network instance is generated by a 2D point process with a given number of nodes $|{\mathcal{V}}|\in\{20,30,\dots,110\}$ uniformly distributed in the plane with a constant density of $8/\pi$. We consider a simplified scenario where all mobile devices are equipped with omnidirectional antennas and transmit at identical power levels, which can be captured by the unit-disk interference model previewed in Section III-A. A link is established between two nodes if they are located within a distance of $1$ unit. The average degree of conflict graphs under this test configuration is $34.6$. For each network size $|{\mathcal{V}}|$, we generate 100 test instances by drawing 10 instances of random networks, each with 10 realizations of random source-destination pairs and random link rates. A test instance contains a number (uniformly chosen between $\lfloor 0.30|{\mathcal{V}}|\rfloor$ and $\lceil 0.50|{\mathcal{V}}|\rceil$) of random flows between different pairs of sources and destinations. Each test instance also includes a realization of uniformly distributed long-term link rates, $r_{e}\sim\mathbb{U}(10,42)$, which are the expected values of the real-time link rates, ${\mathbf{R}}_{e,t}\sim\mathbb{N}(r_{e},3)$, truncated to between $r_{e}\pm 9$. The link rates are configured to capture fading channels with lognormal shadowing. A test instance is simulated for a total of $T=1000$ time steps.
@@
Two types of flows are employed: in a streaming flow $f$, new packets arrive at the source following a Poisson process with a uniformly random arrival rate $\lambda(f)\sim\mathbb{U}(0.2,1.0)$ for all time slots; in a bursty flow, packets follow Poisson arrivals with a random rate $\lambda(f)\sim\mathbb{U}({2.0,10.0})$ (lighter traffic) or $\lambda(f)\sim\mathbb{U}({6.6,33.0})$ (heavier traffic) for $t<30$, and no new packet arrivals for $t\geq 30$. With $T=1000$, a streaming and a heavier bursty flows are expected to inject similar total numbers of packets into the network, allowing better assessment of the impact of traffic pattern. These flow rates are considered lightweight since the expected exogenous packet arrival rate per node of $0.24$ is only $8\%$ of the expected total maximum outflow rate per node of $3$, see Appendix B for details. The end-to-end delay of a tested BP scheme is collected by tracking the time each packet arrived at the source node $t_{0}$ and the destination node (departure time) $t_{1}$. The end-to-end delay of an undelivered packet is treated as $T$. We record the latency of individual flows (source-destination pairs) by averaging the end-to-end delays of their packets.
TABLE I: Hyperparameters for Training
@@
GCNN config $L=5$, $g_{l}=32$ for $21\leq l\leq 4,g_{0}=1,g_{L}=$ 2, see Section IV-A for other details
@@
TABLE II: Tested low-overhead BP algorithms and their detailed configurations, information requirements, and result summary.
@@
Acronym Distance Scaling $U_{i}^{(c)}(t)$ Info. Info. latency (s) delivery (s) latency (b) delivery (b) network capacity (s)
@@
$t_{0}$: the time a packet arrives to the network, $r$: individual long-term link rate, $\bar{r}$: network average long-term link rate, $x$: predicted link duty cycle, $Q$: queue length, SJB: total sojourn time of all packets in a queue, HOL: sojourn time of head-of-line packet in a queue, expQ: backlog function defined in (11).
@@
We train a 5-layer GCNN ($L=5$, $g_{l}=32,l\in\{1,\dots,4\}$) as detailed in Section IV-A according to the RL scheme in Section IV-D. The training instances are generated similar to the test instances described earlier, except that the underlying network instances are from a set of 100 random networks with $|{\mathcal{V}}|\in\{20,30,\dots,60\}$, and the number of flows is uniformly chosen between $\lfloor 0.15|{\mathcal{V}}|\rfloor$ and $\lceil 0.30|{\mathcal{V}}|\rceil$. Notice that our test instances cover larger networks and denser traffic than our training configuration, to illustrate that our GNN-based SP-BP schemes can generalize to problems of larger scales. The detailed hyperparameters and training setting are listed in Table I, and the model converged after 5 epochs. [^1] Training takes 5 hours on a workstation with a specification of 32GB memory, 8 cores, and Geforce GTX 1070 GPU. The source code is published at https://github.com/zhongyuanzhao/dutyBP
@@
The tested BP algorithms (acronyms in parentheses identify legends in figures) include 1) unbiased BP: basic BP (BP), delay-based unbiased BP with SJB [21] (BP-SJB) and HOL [19] (BP-HOL); 2) queue-length based biased BP: EDR [12] with edge weight $\delta_{e}=\bar{r}$ (EDR-$\bar{r}$), and GCN-based delay-aware shortest path bias [1] (SP-$\bar{r}/(xr)$) with edge weight $\delta_{e}=\bar{r}/(x_{e}r_{e})$, and its scaled version (SP-$\bar{r}/(xr)$-min); and 3) delay-based biased BP: combination of BP-SJB and EDR-$\bar{r}$ (EDR-$\bar{r}$-SJB), combination of BP-HOL and EDR-$\bar{r}$ (EDR-$\bar{r}$-HOL), and the two biased BP schemes based on expQ (EDR-$\bar{r}$-expQ and SP-$\bar{r}/(xr)$-expQ) where $\epsilon=0.01$ in (11). Postfix -$\min$ in legend refers to a SP-BP scheme with optimally scaled edge weight $\tilde{\delta}_{e}={\delta}_{e}\bar{r}/(\min_{e\in{\mathcal{E}}}{\delta}_{e})$. The details of these BP schemes, including their requirements on local and global information and test rankings, are listed in Table II.
VII-A Bias Scaling
We first examine the optimal bias scaling approach proposed in Section V-A. We test EDR-$\delta$ and SP-$\bar{r}/(xr)$-min with edge weight adjustment $\underset{e\in{\mathcal{E}}}{\min}\;\delta_{e}\vcentcolon=a\bar{r}$, for $a\in\left[0.5,1.5\right]$, on random networks of 100 nodes. The results in Fig. 4 show that $a=1.0$ is indeed the optimal setting for EDR-$\delta$ and near-optimal setting for SP-$\bar{r}/(xr)$-min, under both streaming and bursty traffic. This provides empirical validation of the adjustment rule $\underset{e\in{\mathcal{E}}}{\min}\;\delta_{e}\vcentcolon=\bar{r}$ advocated in Section V. Notice that the settings tested in Fig. 4 go beyond the simplified setting of CFLP with constant rates treated in Section V. However, the proposed scaling is still empirically optimal in these broader scenarios. Since SP-$\bar{r}/(xr)$-min can tolerate a wide range of $a$, to keep its distributed execution, it can be implemented based on statistical information of edge weights rather than their global minimum. The different sensitivities of the end-to-end delay to the scaling of edge weight under the two tested SP-BP schemes, as shown in Fig. 4, reveal that the shortest path based on link features [1] not only achieves better performance but is also more robust to the scaling choice.
VII-B Network Size and Traffic Type

### v2

@@ Abstract
{IEEEkeywords}
Backpressure routing, MaxWeight scheduling, shortest path, queueing networks, last packet problem, graph neural networks.
1 INTRODUCTION
@@
There are two categories of low-overhead BP variations for latency improvement: 1) Biased BP adds pre-defined queue-agnostic biases, e.g., (functions of) shortest path distance [12, 13, 15, 1, 2] to the backlog metric. Shortest path-biased BP (SP-BP) can mitigate the drawbacks of slow startup and random walk, while maintaining throughput optimality [12, 13, 15] at a low cost, i.e., a one-time communication overhead for bias computation. As exemplified in Fig. 1(b), two major routes around the empty central area can be quickly established by SP-BP [12, 13]. 2) Delay-based BP employ delay-based backlog metrics rather than queue length [19, 16, 21], in order to address the last packet problem and to stabilize the queues in the network. Importantly, the aforementioned approaches do not require additional communication overhead at every time step, unlike those based on queue state information (QSI) of the network [16, 17] or virtual queues [20, 18] (see Section 2.1 for a more detailed survey). However, to get the most out of these approaches, careful parameter tuning is required, typically done via trial-and-error.
@@
2 RELATED WORK
2.1 BACKPRESSURE ROUTING
@@
2.2 GNNS FOR WIRELESS RESOURCE ALLOCATION
@@
Figure 2: System diagram of our GNN-enhanced SP-BP scheme.
3 SYSTEM MODEL AND BACKPRESSURE ALGORITHM
@@ 3.2 BACKPRESSURE ALGORITHM
where $U_{i}^{(c)}(t)$ is the backlog metric, whose relationship with queue lengths is detailed in Section 3.3. For shorter notation, we define the backpressure of commodity $c$ on the directed link ($\overrightarrow{i,j}$) as $U_{ij}^{(c)}(t)=U_{i}^{(c)}(t)-U_{j}^{(c)}(t)$. In step 2, the maximum backpressure of ($\overrightarrow{i,j}$) is found as:
@@
3.3 (BIASED) BACKLOG METRICS
@@
4 LINK DUTY CYCLE PREDICTION WITH GNNS
Existing SP-BP algorithms are based on the hop distance (minimal hop count) between two nodes in the networks, which implicitly assumes that every link contributes equally to the end-to-end delay of routing. This, however, ignores the fact that the average delay of each link is different due to its network locality and link scheduling policy. In MaxWeight scheduling, the likelihood of a link being scheduled depends on its local conflict or interference topology and the network traffic load. We propose using a graph convolutional neural network (GCNN) to predict the expected delay on each link, which can serve to design a shortest path bias between node $i$ and the destination node $c$ in (6) better than the shortest hop distance. The GNN and its training setup are illustrated in the upper right blocks of Fig. 2, while the optimal scaling of link weights is discussed in Section 5.1. Since our focus is on model-based framework that enables the applications of GNNs for enhancing SP-BP, we select GCNN for its simplicity and good generalizability, leaving the search of optimal GNN architectures and hyperparameters for the future work.
4.1 GRAPH CONVOLUTIONAL NEURAL NETWORKS
@@
4.2 DELAY-AWARE EDGE WEIGHTS
Depending on the availability of the long-term link rate $r_{e}$ and link duty cycle $x_{e}$ for any link $e\in{\mathcal{E}}$, we consider the weight of each link for finding the shortest path distance between any pair of nodes to be defined in three possible ways: $\delta_{e}=1/x_{e}$, $\delta_{e}=1/r_{e}$, or $\delta_{e}=\bar{r}/(x_{e}r_{e})$, where $\bar{r}$ is the network-wide average link rate defined in Section 3.1. In addition, these link weights can be optimally scaled by $\bar{r}/\delta_{\min}$ based on the principle in Section 5.1. The computation of vector $\bm{\delta}=\left[\delta_{e}|e\in{\mathcal{E}}\right]$ is denoted as a function $\bm{\delta}=f_{w}({\mathbf{x}},{\mathbf{r}},\bar{r})$. By setting the edge weights of the connectivity graph ${\mathcal{G}}^{n}$ as $\bm{\delta}$, bias $B_{i}^{(c)}$ is set as the weighted shortest path distance between nodes $i$ and $c$ on ${\mathcal{G}}^{n}$. This distance can be computed with distributed algorithms, such as weighted APSP in general or weighted SSSP when a new node joins the network.
4.3 COMPLEXITY
@@ 4.4 TRAINING
Intuitively, by minimizing the loss in (9), we are choosing parameters $\bm{\omega}$ such that the softmax output of our GCNN ${\mathbf{X}}^{L}$ is close to predicting the fraction of time that each link is scheduled. We do not seek to learn this complex function for any specific topology, but rather, we want to minimize the average error over instances drawn from $\Omega$. In practice, we approximate the expected values in (9) with the corresponding empirical averages. Indeed, with the collected experience tuples, we update the parameters $\bm{\omega}$ of the GCNN after each training instance, through batch training with random memory sampling, employing the Adam optimizer. Notice that, once trained, the GCNN can be used to compute the delay-aware bias for previously unseen topologies without any retraining. As long as the new topology, arrival rates, and link rates are similar to those observed in $\Omega$ during training, we illustrate in Section 7 that the trained GCNN generalizes to new test instances not observed during training.
5 IMPROVEMENTS TO BIASED BACKPRESSURE
@@
5.1 OPTIMAL SCALING OF EDGE WEIGHTS
To the best of our knowledge, the constant scale of the edge weights $\delta_{e},e\in{\mathcal{E}}$ for computing the shortest path biases in SP-BP schemes are often selected via trail-and-error, such as a constant $\delta_{e}$ for all $e\in{\mathcal{E}}$ in EDR [12] and $\delta_{e}=\bar{r}/(x_{e}r_{e})$ proposed in Section 4.2. Our goal is to provide a principled method of optimization for such scaling parameter(s).
@@ 5.1 OPTIMAL SCALING OF EDGE WEIGHTS
Proof 5.1.
In queue length-based SP-BP, the backpressure on link $(\overrightarrow{i,j})$ is $U_{ij}^{(c)}\!(t)=Q_{ij}^{(c)}\!(t)+\big{(}B_{i}^{(c)}\!-\!B_{j}^{(c)}\big{)}=q+\delta_{e}$. Since link $(\overrightarrow{i,j})$ is scheduled at time $t$, we have that $U_{ij}^{(c)}\!(t+1)=\delta_{e}-q$. To avoid the direction of the backpressure on link $(i,j)$ being reversed at time $t+1$, i.e., $U_{ij}^{(c)}\!(t\!+\!1)<0$, we need to set $\delta_{e}\geq\bar{r}$.
@@
Although the above discussion is based on the CFLP setting with all link rates equal to the average $\bar{r}$, in Section 7, we demonstrate experimentally the optimality of the above choice under heterogeneous link rates and general traffic settings.
5.2 BIAS MAINTENANCE
@@
5.3 DELAY-AWARE BACKLOG METRIC: EXPQ
@@
where $\epsilon\geq 0$ is a small constant parameter, and it follows that $g\big{(}{\mathcal{Q}}_{i}^{(c)}(t)\big{)}=Q_{i}^{(c)}(t)$ for $\epsilon=0$ and that $g\left(\varnothing\right)=0$ for an empty queue $\varnothing$. With this definition, it is easy to find that $g\big{(}{\mathcal{Q}}_{i}^{(c)}(t)\big{)}\geq Q_{i}^{(c)}(t)$. Further restrictions on $\epsilon$ can be found in (25) and the following discussion in Section 6.2.
@@ 6 THROUGHPUT OPTIMALITY
6.1 QUEUE LENGTH-BASED SP-BP SCHEME
@@
By squaring the queue update in (15), and using the fact that $(\max[q-o,b]+p)^{2}\leq q^{2}+o^{2}+p^{2}+b^{2}+2bp+2q(p-o)$ for $q\geq 0,o\geq 0,b\geq 0,p\geq 0$ (see Appendix 9), we have:
@@
Next, we show that the biased BP described in Section 3.2 is designed to choose a feasible policy ${\mathbf{M}}(t)=\big{(}\mu_{ij}^{(c)}(t)\big{)}_{i,j,c\in{\mathcal{V}}}$ under current network state to minimize the bound of the Lyapunov drift, i.e., the right-hand side of (16). Since $W$ and $\lambda_{i}^{(c)}$ are constants, this amounts to minimizing the following:
@@
Since $W\!<\!\infty$, $0\!\leq\!B_{i}^{(c)}\!<\!\infty$, the condition of strong stability in (12) immediately follows by moving $\sum_{i,c\in{\mathcal{V}}}B_{i}^{(c)}$ from the left hand side of (21) to the right hand side.
6.2 SP-BP SCHEME WITH EXPQ
Theorem 6.4.
There exists a constant $\epsilon\geq 0$ such that SP-BP with delay-based backlog metric expQ defined in Section 5.3, denoted as $E_{i}^{(c)}\!(t)=g\!\left(\!{\mathcal{Q}}_{i}^{(c)}(t)\!\right)$, meets the condition of strong stability in (12) within a reduced network capacity region.
Proof 6.5.
@@ 6.2 SP-BP SCHEME WITH EXPQ
We apply the proof for Theorem 6.2 by substituting (15) with (22), and replacing $U_{i}^{(c)}(t)$ by $\tilde{U}_{i}^{(c)}(t)$ and $A_{i}^{(c)}\!(t)$ by $A_{i}^{(c)}\!(t)+b$, thus obtain the policy of SP-BP with expQ as
@@
With (24), (25) and ${E}_{i}^{(c)}\!(t)\geq{Q}_{i}^{(c)}\!(t)$, all queues are strongly stable within the network capacity region shrank by $b$.
@@
Simulations in [13, Ch 4.9] and Section 7.3 show that for BP with queue length-based backlog metrics, the inclusion of non-negative queue-agnostic biases does not shrink the network capacity region of the basic BP. However, as illustrated in Section 7.3, the application of delay-based backlog metrics, including expQ, will reduce the network capacity of BP.
7 NUMERICAL EXPERIMENTS
@@
To illustrate the effectiveness of our proposed enhancements to the SP-BP schemes in improving routing performance, we evaluate a series of low-overhead BP routing algorithms in simulated wireless multi-hop networks. Each network instance is generated by a 2D point process with a given number of nodes $|{\mathcal{V}}|\in\{20,30,\dots,110\}$ uniformly distributed in the plane with a constant density of $8/\pi$. We consider a simplified scenario where all mobile devices are equipped with omnidirectional antennas and transmit at identical power levels, which can be captured by the unit-disk interference model previewed in Section 3.1. A link is established between two nodes if they are located within a distance of $1$ unit. The average degree of conflict graphs under this test configuration is $34.6$. For each network size $|{\mathcal{V}}|$, we generate 100 test instances by drawing 10 instances of random networks, each with 10 realizations of random source-destination pairs and random link rates. A test instance contains a number (uniformly chosen between $\lfloor 0.30|{\mathcal{V}}|\rfloor$ and $\lceil 0.50|{\mathcal{V}}|\rceil$) of random flows between different pairs of sources and destinations. Each test instance also includes a realization of uniformly distributed long-term link rates, $r_{e}\sim\mathbb{U}(10,42)$, which are the expected values of the real-time link rates, ${\mathbf{R}}_{e,t}\sim\mathbb{N}(r_{e},3)$, truncated to between $r_{e}\pm 9$. The link rates are configured to capture fading channels with lognormal shadowing. A test instance is simulated for a total of $T=1000$ time steps.
@@
Two types of flows are employed: in a streaming flow $f$, new packets arrive at the source following a Poisson process with a uniformly random arrival rate $\lambda(f)\sim\mathbb{U}(0.2,1.0)$ for all time slots; in a bursty flow, packets follow Poisson arrivals with a random rate $\lambda(f)\sim\mathbb{U}({2.0,10.0})$ (lighter traffic) or $\lambda(f)\sim\mathbb{U}({6.6,33.0})$ (heavier traffic) for $t<30$, and no new packet arrivals for $t\geq 30$. With $T=1000$, a streaming and a heavier bursty flows are expected to inject similar total numbers of packets into the network, allowing better assessment of the impact of traffic pattern. These flow rates are considered lightweight since the expected exogenous packet arrival rate per node of $0.24$ is only $8\%$ of the expected total maximum outflow rate per node of $3$, see Appendix 10 for details. The end-to-end delay of a tested BP scheme is collected by tracking the time each packet arrived at the source node $t_{0}$ and the destination node (departure time) $t_{1}$. The end-to-end delay of an undelivered packet is treated as $T$. We record the latency of individual flows (source-destination pairs) by averaging the end-to-end delays of their packets.
Table 1: Hyperparameters for Training
@@
GCNN config $L=5$, $g_{l}=32$ for $21\leq l\leq 4,g_{0}=1,g_{L}=$ 2, see Section 4.1 for other details
@@
Table 2: Tested low-overhead BP algorithms and their detailed configurations, information requirements, and result summary.
@@
Acronym Distance Scaling $U_{i}^{(c)}(t)$ Info. Info. latency (s) delivery (s) latency (b) delivery (b) network cap.(s)
@@
$t_{0}$: the time a packet arrives to the network, $r$: individual long-term link rate, $\bar{r}$: network average long-term link rate, $x$: predicted link duty cycle, $Q$: queue length, SJB: total sojourn time of all packets in a queue, HOL: sojourn time of head-of-line packet in a queue, expQ: backlog func. defined in (11).
@@
We train a 5-layer GCNN ($L=5$, $g_{l}=32,l\in\{1,\dots,4\}$) as detailed in Section 4.1 according to the RL scheme in Section 4.4. The training instances are generated similar to the test instances described earlier, except that the underlying network instances are from a set of 100 random networks with $|{\mathcal{V}}|\in\{20,30,\dots,60\}$, and the number of flows is uniformly chosen between $\lfloor 0.15|{\mathcal{V}}|\rfloor$ and $\lceil 0.30|{\mathcal{V}}|\rceil$. Notice that our test instances cover larger networks and denser traffic than our training configuration, to illustrate that our GNN-based SP-BP schemes can generalize to problems of larger scales. The detailed hyperparameters and training setting are listed in Table 1, and the model converged after 5 epochs. [^1] Training takes 5 hours on a workstation with a specification of 32GB memory, 8 cores, and Geforce GTX 1070 GPU.
@@
The tested BP algorithms (acronyms in parentheses identify legends in figures) include 1) unbiased BP: basic BP (BP), delay-based unbiased BP with SJB [21] (BP-SJB) and HOL [19] (BP-HOL); 2) queue-length based biased BP: EDR [12] with edge weight $\delta_{e}=\bar{r}$ (EDR-$\bar{r}$), and GCN-based delay-aware shortest path bias [1] (SP-$\bar{r}/(xr)$) with edge weight $\delta_{e}=\bar{r}/(x_{e}r_{e})$, and its scaled version (SP-$\bar{r}/(xr)$-min); and 3) delay-based biased BP: combination of BP-SJB and EDR-$\bar{r}$ (EDR-$\bar{r}$-SJB), combination of BP-HOL and EDR-$\bar{r}$ (EDR-$\bar{r}$-HOL), and the two biased BP schemes based on expQ (EDR-$\bar{r}$-expQ and SP-$\bar{r}/(xr)$-expQ) where $\epsilon=0.01$ in (11). Postfix -$\min$ in legend refers to a SP-BP scheme with optimally scaled edge weight $\tilde{\delta}_{e}={\delta}_{e}\bar{r}/(\min_{e\in{\mathcal{E}}}{\delta}_{e})$. The details of these BP schemes, including their requirements on local and global information and test rankings, are listed in Table 2.
@@
7.1 BIAS SCALING
We first examine the optimal bias scaling approach proposed in Section 5.1. We test EDR-$\delta$ and SP-$\bar{r}/(xr)$-min with edge weight adjustment $\underset{e\in{\mathcal{E}}}{\min}\;\delta_{e}\vcentcolon=a\bar{r}$, for $a\in\left[0.5,1.5\right]$, on random networks of 100 nodes. The results in Fig. 4 show that $a=1.0$ is indeed the optimal setting for EDR-$\delta$ and near-optimal setting for SP-$\bar{r}/(xr)$-min, under both streaming and bursty traffic. This provides empirical validation of the adjustment rule $\underset{e\in{\mathcal{E}}}{\min}\;\delta_{e}\vcentcolon=\bar{r}$ advocated in Section 5. Notice that the settings tested in Fig. 4 go beyond the simplified setting of CFLP with constant rates treated in Section 5. However, the proposed scaling is still empirically optimal in these broader scenarios. Since SP-$\bar{r}/(xr)$-min can tolerate a wide range of $a$, to keep its distributed execution, it can be implemented based on statistical information of edge weights rather than their global minimum. The different sensitivities of the end-to-end delay to the scaling of edge weight under the two tested SP-BP schemes, as shown in Fig. 4, reveal that the shortest path based on link features [1] not only achieves better performance but is also more robust to the scaling choice.
7.2 NETWORK SIZE AND TRAFFIC TYPE